OpenStudy (anonymous):

PLEASE HELP!!!:) Combine like terms. 7/8x - 1/2x A.1/8x B.3/8 C. 3/8x D. x

OpenStudy (firejay5):

this is a math question not an English question; answer is: C. 3/8x, because they have the same variable, so you can subtract them 7/8 = 0.875, 1/2 = 0.50 <--- subtract 0.875 from 0.50 and you get 0.375. A and D isn't correct, so it's either B or C; well it's C. since they have the same variable (x)

OpenStudy (anonymous):

is the x in the denominator?

OpenStudy (anonymous):

or in the numerator?

OpenStudy (firejay5):

It can be written in two ways: \[\frac{ 1 }{ 8x } or \frac{ 1 }{ 8 }x\]

OpenStudy (anonymous):